The following data for a hypothetical country in millions of dollars for the year 2020. Depreciation =200, Exports =150, public transfer =200, Gross domestic private investment =300, corporate income tax =100, factor receipts from abroad =400, gevenment expenditure on goods and services =250, interest income =800, compensation of employees =2600, net interest on government debit =50, indirect business taxes =100, factor payments to abroad =200, imports =200, proprietors incom(profits) =700, retained corporate profit =200, personal consumption expenditure =4500, personal taxes =100, social security contribution =50, rental income =600. A. Calculate the gross domestic product of the country I. Using the expenditure approach II. Using the income approach B, Calculate the gross nationalproduct of the country. C, calculate the real GDP of the nation if the price index for the year is 111. D, calculate the nations growth rate lf the real GDP for the year 2019 is 4000.
1
Expert's answer
2021-09-02T10:47:18-0400

a)

i)GDP @MP calculation by expenditure method (2020)



Consumption Expenditure = Personal consumption expenditure+ Personal taxes+ Contribution to social security

                        "= 4500+ 100+ 250 = 4850"

Investment Expenditure   = 300

Government Expenditure  = Govt. Expenditure - Interest on Government Debit "= 250- 50 = 200"

Net export "= Export - Import = 150-200 = -50."


ii) GDP@MP calculation by income method (2020)



Profit = Proprietors Income + Retain Corporate Profit + Corporate Income Tax = 700+200+100 = 1000 


b)

  • Gross National Product at market price of the country = GDP @MP of the country + Factors receipts from abroad- Factors receipts to abroad

                                                 "= 5300 + 400 - 200\\

 = \$ 5500 Million"


c)

  • Real GDP of the Nation  "=\frac{ (Nominal \space GDP X 100) }{ 111}"

                                             


"= \frac{ (5300 X 100)}{ 111}\\

 = \$ 4774.77 Million"

d)

Nation's Growth Rate"= \frac{{GDP(2020) - GDP (2019)}\times100 }{ GDP(2019}"

                                                "= \frac{(4774.77-4000)\times100}{ 4000}\\

 = \frac{774.77\times100}{4000}\\

 = 19.37 %"